# ROOFS: RObust biOmarker Feature Selection

### v1

@@ 4 Discussion
This paper introduced roofs, a Python framework designed to help researchers identify an FS method adapted to their predictive task. FS is known to be data-dependent, with method performance varying across datasets and no single approach being universally optimal [42, 25, 14]. However, FS methods are often adopted from the literature or research teams established practices, with little consideration of their suitability for the specific dataset at hand. roofs addresses this gap by integrating the evaluation of over 30 FS methods from different algorithmic families into a single pipeline that requires low to zero additional programming. It generates automatic reports summarizing FS performance metrics, as well as measures of individual feature robustness. A key element of roofs is its emphasis on stability – a property of FS that is frequently ignored in biomedical studies and in the development of new FS approaches despite its importance for reproducibility.
To our knowledge, no previously released packages combines multiple pre-implemented FS methods with the optimism correction framework for model evaluation. The closest existing tool is the R package pminternal (https://github.com/stephenrho/pminternal), which estimates optimism for user-defined model functions, including those that contain an FS step. However, it does not provide any pre-implemented FS methods.
@@
We demonstrated the utility of roofs through application to the dataset from the PIONeeR clinical trial, focused on predicting resistance to anti-PD-(L)1 immunotherapy in advanced lung cancer. It enabled identification of the adjusted p-values filter – the union of a t-test and logistic regression controlled for PD-L1 – as an FS method offering the best stability, confidence in selected features, predictive performance and balance between TPR and FPR. The general findings of this study align with previous benchmarks where simple statistical filters outperformed more complicated ML approaches [20]. Therefore, we recommend starting the analysis of new datasets by benchmarking these methods, given their additional advantage of computational efficiency.
To conclude, the choice of an FS method and, therefore, the resulting subset of features or "signature", should be driven by the dataset’s characteristics and study goals. It should consider study-specific acceptable trade-off between model stability, performance, and relative costs of false positives and false negatives. roofs can serve as a practical tool for researchers working in settings where reproducibility is crucial. By enabling comprehensive benchmarking, roofs and similar tools have the potential to improve the robustness and reproducibility of biomarker discovery and increase the translational value of clinical predictive models.
@@
This work benefited from a government grant handled by the French National Research Agency (ANR) as part of the France 2030 investment plan, under the reference ANR-17-RHUS-0007.
@@
The authors gratefully acknowledge the support of the APHM, which sponsored the PIONeeR clinical studies. Its role was to control the appropriateness of ethical and legal considerations for all centers and to perform the monitoring of the consents signed and the clinical data recorded and coded as part of the study. The authors are grateful to all the patients and their families, as well as all the investigators, for their participation in the study.
This work was also supported by a grant from the French government, managed by the National Research Agency (ANR), under the France 2030 program within the DIGPHAT project, reference ANR-22-PESN-0017.
@@
Bolasso Modified bootstrap LASSO. Selects features with a bootstrap frequency of selection of $\geq 0.5$ instead of 1. [1]
@@
Block HSIC LASSO Selects features based on their dependence with the outcome using the Hilbert-Schmidt Independence Criterion. [11]
@@
Stability selection LASSO-based subsampling method that performs repeated fits across a range of regularization parameters and selects features selected with high frequency above a specified threshold. [34]
Stabl LASSO Recent extension of stability selection that injects artificial noise variables into the data to enable data-driven choice of the selection frequency threshold, with the goal of minimizing false discoveries. [21]
@@
ReliefF Ranks features by a distance-based score and selects top-k features. For each sample, the algorithm penalizes the features with different values in the 5 nearest neighbors of the same class and rewards the features with different values for the 5 nearest neighbors of the opposite class. [31, 40]
@@
Hierarchical clustering Performs hierarchical clustering of features based on Spearman correlation coefficients, with the number of clusters equal to the desired number of features to be selected. From each cluster, the feature with the highest point biserial correlation with the outcome is selected.
p.adjust t-test + logistic .05 Union-based filter that combines features selected by t-test and logistic regression (controlling for user-specified covariates; PD-L1 in the case of PIONeeR data), where p-values from each test were adjusted using the Benjamini-Hochberg procedure with a 0.05 threshold.
@@
LR-RFE Recursive feature elimination that iteratively removes 10% of features with the lowest coefficients from a logistic regression model with the L2 penalty until the pre-specified number of features is reached. [18]
@@
Random FS Selects a feature subset randomly. For the benchmark on PIONeeR data, subset sizes $k$ were randomly sampled from the range 1-20 for the real dataset and set to $k=19$ for the semi-synthetic dataset, corresponding to the number of predictors in the true model.

### v2

@@ 4 Discussion
This paper introduced ROOFS, a Python framework designed to help researchers identify an FS method adapted to their predictive task. FS is a methodologically mature field, with extensive advances in algorithmic development [34], stability assessment [41], and false discovery control [24]. However, many of the theoretical developments have not been fully adopted in applied biomedical research, where practitioners often rely on a relatively narrow set of familiar methods. Moreover, FS is known to be data-dependent, with method performance varying across datasets and no single approach being universally optimal [28, 47, 16]. Nevertheless, FS methods are often applied with little consideration of their suitability for the specific dataset at hand. ROOFS addresses these gaps by integrating the evaluation of over 30 FS methods from different algorithmic families into a single pipeline that requires low to zero additional programming. To facilitate method comparison, ROOFS generates automatic reports summarizing FS performance metrics, as well as measures of individual feature robustness. A key element of ROOFS is its emphasis on stability – a property of FS that is frequently ignored in biomedical studies and in the development of new FS approaches despite its importance for reproducibility.
To our knowledge, no previously released packages combine multiple pre-implemented FS methods with the optimism correction framework for model evaluation. The closest existing tool is the R package pminternal (https://github.com/stephenrho/pminternal), which estimates optimism of the user-defined model functions, including those that contain an FS step. However, it does not provide any pre-implemented FS methods.
@@
We demonstrated the utility of ROOFS through application to the dataset from the PIONeeR clinical trial, focused on predicting resistance to anti-PD-(L)1 immunotherapy in advanced lung cancer. It enabled identification of the adjusted p-values filter – the union of a t-test and logistic regression controlled for PD-L1 at a 0.05 threshold – as an FS method offering the best stability, confidence in selected features, predictive performance and balance between TPR and FPR. The general findings of this study align with previous benchmarks where simple statistical filters outperformed more complicated ML approaches [23]. Therefore, we recommend starting the analysis of new datasets by benchmarking these methods, given their additional advantage of computational efficiency. ROOFS was also recently successfully applied to the analysis of the SChISM clinical study (Size CfDNA Immunotherapy Signature Monitoring, NCT05083494). It allowed to identify a nine-variables signature derived from fragmentomics circulating free DNA data, with enhanced predictive power of resistance (AUC = 0.89 ± 0.033, PPV = 0.69 ± 0.075) [40].
To conclude, the choice of an FS method and, therefore, the resulting subset of features or "signature", should be driven by the dataset’s characteristics and study goals. It should consider study-specific acceptable trade-off between model stability, performance, and relative costs of false positives and false negatives. ROOFS can serve as a practical tool for researchers working in settings where reproducibility is crucial. By enabling comprehensive benchmarking, ROOFS and similar tools have the potential to improve the reproducibility of biomarker discoveries and increase the translational value of clinical predictive models.
Key points
  • There is a significant disconnect between the rich methodological development of feature selection (FS) and its use in applied biomedical research. Practitioners frequently rely on a limited subset of FS approaches and do not use comprehensive evaluation procedures to assess the suitability of the method for a given dataset.
  • FS stability, defined as the consistency of selected features across data resamples, is essential for reproducibility and clinical translation of biomarker signatures, yet it often remains unreported in biomedical studies.
  • We developed a Python package, ROOFS, designed to benchmark multiple FS methods on user-provided data with emphasis on method stability, individual feature robustness, predictive performance, and error rates in a separate analysis with a simulated outcome.
  • We demonstrated ROOFS utility by applying it to a real-world problem: predicting resistance to immunotherapy in lung cancer patients from the PIONeeR clinical trial. This comparative study identified a statistical filter based on the union of Benjamini-Hochberg adjusted p-values from t-test and logistic regression (adjusted for PD-L1) at a 0.05 threshold as the optimal FS method across multiple evaluation metrics.
Data availability statement
The PIONeeR data underlying this article cannot be shared publicly due to patient privacy reasons. Two publicly available datasets are provided in the package for tutorial and testing [9, 39, 40], as well as one synthetic dataset.
@@
This work was supported by the French National Research Agency (ANR) under the France 2030 investment plan through grants ANR-17-RHUS-0007 and ANR-22-PESN-0017, the latter awarded within the DIGPHAT project.
@@
Acknowledgements
The authors gratefully acknowledge the support of the APHM, which sponsored the PIONeeR clinical study. Its role was to control the appropriateness of ethical and legal considerations for all centers and to perform the monitoring of the consents signed and the clinical data recorded and coded as part of the study. The authors are grateful to all the patients and their families, as well as all the investigators, for their participation in the study.
The authors are grateful to Linh Nguyen Phuong, PhD, and Mohamed Boussena for helpful comments.
@@ Acknowledgements
Bolasso Modified bootstrap LASSO. Selects features with a bootstrap frequency of selection of $\geq 0.5$ instead of 1. Our modification of [1]
Stability selection LASSO-based subsampling method that performs repeated fits across a range of regularization parameters and selects features selected with a frequency above a specified threshold. [37]
Stabl LASSO Recent extension of stability selection that injects artificial noise variables to enable data-driven choice of the selection frequency threshold. [24]
@@
Block HSIC LASSO Method that the Hilbert-Schmidt Independence Criterion (HSIC) to identify non-linear dependencies between features and the outcome while penalizing redundant features. To reduce memory complexity for the case of high-dimensional data, the method partitions samples into smaller blocks. The final bagging block HSIC estimator is computed by averaging HSIC values across multiple random permutations of the data partitions. [13]
@@
ReliefF Ranks features by a distance-based score and selects top-k features. For each sample, the algorithm penalizes the features with different values in the nearest neighbors (5 by default) of the same class and rewards the features with different values for the nearest neighbors (similarly, 5 by default) of the opposite class. [34, 46]
@@
Hierarchical clustering Performs hierarchical clustering of features based on Spearman correlation coefficients, with the number of clusters equal to the desired number of features to be selected. From each cluster, the feature with the highest point biserial correlation with the outcome is selected. Our modification of [43]
p.adjust t-test + logistic .05 Union-based filter that combines features selected by t-test and logistic regression (controlling for user-specified covariates; PD-L1 in the case of PIONeeR data), where p-values from each test were adjusted using the Benjamini-Hochberg procedure with a 0.05 threshold. Our modification of p-value filtering
@@
LR-RFE Recursive feature elimination that iteratively removes 10% of features with the lowest absolute coefficients from a logistic regression model with the L2 penalty until the pre-specified number of features is reached. [21]
@@
Random FS Selects a feature subset randomly. For the benchmark on PIONeeR data, subset sizes $k$ were randomly sampled from the range 1-20 for the real dataset. For the semi-synthetic dataset, subset size was either set to $k=19$ (a fixed size corresponding to the number of predictors in the true model) or sampled randomly from [1, $p$].